OpenStudy (anonymous):

tan inverse 2x/1+15x^2

OpenStudy (anonymous):

What has to be done here?

OpenStudy (anonymous):

\[\tan^{-1} \frac{2x}{1+15x^2}\]

OpenStudy (anonymous):

differentiate

OpenStudy (anonymous):

Ok pls wait

OpenStudy (anonymous):

Do you know the differentiation of \[\tan^{-1}x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It is \[\tan^{-1}x = \frac {1}{1+x^2}\] so using it we can

OpenStudy (anonymous):

can u solve it if u dont mind

OpenStudy (anonymous):

Therefore let \[y= \tan^{-1} \frac{2x}{1+15x^2} \rightarrow y' = \frac{d}{dx}\tan^{-1} \frac{2x}{1+15x^2}\] \[y' = \frac{1}{1+(\frac{2x}{1+15x^2})^2} \frac{d}{dx}(\frac{2x}{1+15x^2})\] \[y' = \frac{1}{1+\frac{4x^2}{(1+15x^2)^2}} \frac{d}{dx}(\frac{2x}{1+15x^2})\]\[y' = \frac{1}{\frac{(1+15x^2)^2+4x^2}{(1+15x^2)^2}} \frac{d}{dx}(\frac{2x}{1+15x^2})\] \[y' = \frac{(1+15x^2)^2}{(1+15x^2)^2+4x^2} (\frac{(1+15x^2) \frac{d}{dx}2x-2x \frac{d}{dx}(1+15x^2)}{(1+15x^2)^2})\] \[y' = \frac{(1+15x^2)^2}{(1+15x^2)^2+4x^2} (\frac{(1+15x^2) 2-2x (30x)}{(1+15x^2)^2})\] \[y' = (\frac{(1+15x^2) 2-2x (30x)}{(1+15x^2)^2+4x^2})\] \[y' = (\frac{2+30x^2-60x^2}{(1+15x^2)^2+4x^2})\] \[y' = \frac{2-30x^2}{(1+15x^2)^2+4x^2}= \frac{2-30x^2}{1+225x^4+30x^2+4x^2}\] \[y' = \frac{2-30x^2}{1+225x^4+34x^2}\]

OpenStudy (anonymous):

@shalinisengupta

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Medal please

OpenStudy (anonymous):

i m new so dont know how to give medal

OpenStudy (anonymous):

@shalinisengupta